Consider the following data to answer the next 4 questions.

[A] initial (M) [B] initial (M) Initial speed of the reaction (M / s)

0.2 0.1 5

0.2 0.2 10

0.6 0.2 30

3a. The order of the reaction with respect to [A] is__________.

3b. The order of the reaction with respect to [B] is__________.

3c. The net order of the reaction is__________.

3d. The value of k is (include units) is__________.
Answer #1

Rake law :- nate of reaction = K [A]^ [B]Y biven- S.NO [A] [B] speed (role) of reaction - Trail 1 0.2 ol 5 Trail 2 0.2 0.2 10similarly for trail de 10 = K (0.23210.298 – for trail 3 - 30 = K (0.6)? (0.2) 9 - ③ diwidling eq@by eq0 - eq@ = 10K Corofa (dividing eq 0 by eq@ ie - 5 sq - 30 = 106)* congly 10 *10.2) * (oxyg 33con le sue 7 (3) = (3)* o (=) So, take low - I save o